Hydrogen sulfide decomposes according to the following reaction, for whichKc = 9.30 × 10−8 at 700°C: 2 H2S(g) ⇌ 2 H2(g) + S2(g)

If 0.27 mol of H2S is placed in a 3.0−L container, what is the equilibrium concentration of H2(g) at 700°C?

____M
Answer #1

The equilibrium reaction is

2H2S (g) 2H2 ( g) + S2 (g)

Inital concentration of H2S(g) = 0.27/3 = 0.09 M

ICE table for the reaction

[H2S] [ H2] [S2]
Initial 0.09 0 0
Change -2x +2x +x
Equilibrium

(0.09-2x)

2x x

Now, equilibrium constant , Kc = (1)

Putting the values in above equation

Kc = (2x)2×( x) / ( 0.09-2x)2

Given, value of Kc is very small, hence the reaction at equilibrium does not procced to far in the right side , therefore x is very small , therefore we can approximate (0.09-2x) (0.09)

Therefore, Kc = 4x3/(0.09)2

Or, 4x3 = 9.30*10-8×(0.09)2

Or, 4x3 = 0.075×10-8

or, x3 = 1.88×10-10

Or, x = 5.736×10-4 M

Now,

[ S2] = 5.736 ×10-4 M

[ H2S] =2×5.736 ×10-3 M = 11.47×10-4 M

and, [ H2S] = (0.09 -2×5.736×10-4) = 0.0888 M

Now, putting above calculated values in Eq.1

Calculated Kc = (11.47×10-4)2×(5.736×10-4)/(0.0888)2

= 9.56×10-8

Again

x*100÷[ intial concentration of H2S ]

=( 5.736×10-4×100)÷(0.09)

= 0.6376

So, as calculated Kc is nearly equal to given value of Kc and degree of dissociation(x) is less than 5% of initial concentration of H2S, then the approximation was valid.

Therefore, concentration of H2(g) at equilibrium (7000c)

= 11.47×10-4 M

or, 1.147×10-3 M.